# Model Editing at Scale leads to Gradual and Catastrophic Forgetting

### v2

@@ Abstract
Editing knowledge in large language models is an attractive capability that allows us to correct incorrectly learned facts during pre-training, as well as update the model with an ever-growing list of new facts. While existing model editing techniques have shown promise, they are usually evaluated using metrics for reliability, specificity and generalization over one or few edits. We argue that for model editing to have practical utility, we must be able to make multiple edits to the same model. With this in mind, we evaluate current model editing methods at scale, focusing on two state of the art methods - ROME and MEMIT. With the lens of scalability, we evaluate model editing methods for three crucial properties - editing proficiency, fact forgetting and downstream performance. We find that as a model is edited sequentially with multiple facts, it continually becomes less editable, forgets previously edited facts and loses the ability to perform downstream tasks. For ROME and MEMIT, this "forgetting" happens in two phases - an initial gradual but progressive forgetting phase followed by an abrupt or catastrophic forgetting. Both gradual and catastrophic forgetting limit the usefulness of model editing methods at scale - the former makes model editing less effective as multiple edits are made to the model while the latter caps the scalability of such model editing methods. Our analysis also highlights other key limitations of ROME and MEMIT at scale. With our work, we push for better evaluation of model editing and development of model editing methods keeping scalability in mind.
@@ 1 Introduction
zsRE
Who was the architect of Villa Kampen?
CounterFact
Porto is a twin city of
@@ 2 Methods, Models and Datasets
In this paper, we focus on two prominent model editing methods in literature - Rank-One Memory Editing (ROME) Meng et al. (2022a) and Mass-Editing Memory in a Transformer (MEMIT) Meng et al. (2022b), whereas Model Editor Networks using Gradient Decomposition (MEND) Mitchell et al. (2021) and fine-tuning are used as baselines. MEND is a hypernetwork Chauhan et al. (2023) based model editing method that generates the weight updates for the model being edited. ROME and MEMIT first localize knowledge within a model using causal tracing Vig et al. (2020) and then update the weights of the selected layers to inject knowledge. The major difference between ROME and MEMIT is that while ROME works under the assumption that knowledge in LLMs can be updated by updating a single layer, MEMIT updates the weights of multiple layers. We evaluate these model editing methods over two models - GPT2-XL (1.5B) Radford et al. (2019) and GPT-J (6B) Wang and Komatsuzaki (2021). Note that all prior works (Mitchell et al., 2021; Meng et al., 2022a, b) edit base language models and not chat models. We refer the reader to appendix A.1 for a detailed survey of model editing techniques. This section will be added to the main paper upon acceptance.
@@ 3.3 Downstream Evaluation of Edited Models
DISTANCE
$3.339\times 10^{-4}$
$8.156\times 10^{-7}$
@@ Appendix A Appendix
A.1 Related Work
In this paper, we focus on model editing methods that modify the base language model’s parameters. Some of these methods (De Cao et al., 2021; Mitchell et al., 2021) require training a hypernetwork Chauhan et al. (2023) that generates new weights for the model being edited. Other methods (Meng et al., 2022a, b; Li et al., 2023a) directly update specific parts of the model after locating stored facts inside it (Geva et al., 2020). Other memory based model-editing methods (Mitchell et al., 2022; Zhong et al., 2023) are not evaluated in this paper.
A.2 zsRE Compatibility
@@ A.1 Related Work
Task
Few-Shot Prompt
SST-2
Review : excruciatingly unfunny and pitifully unromantic
@@
A.3 Model Editing Implementation Details
@@
A.4 Downstream Evaluation Details
@@
Task
Few-Shot Prompt
MRPC
Are the sentences paraphrases of each other.
@@
Task
Few-Shot Prompt
COLA
Is this sentence linguistically acceptable?
@@
Task
Few-Shot Prompt
NLI
Cyrus captured Babylon without a battle, and remedied the evils done by previous Assyrian and Babylonian rulers by sending prisoners in Babylonia back to their original homelands and aiding in the restoration of temples of the gods of various nations.
@@
A.5 Additional Scaling Experiments
A.5.1 Editing Proficiency
@@ A.6 Generation Examples
Prompt
Generated Text
@@
Table 8: Text generated by GPT-J post the point of catasrophic forgetting when edited using ROME.

### v3

@@ Abstract
Editing knowledge in large language models is an attractive capability that allows us to correct incorrectly learned facts during pre-training, as well as update the model with an ever-growing list of new facts. While existing model editing techniques have shown promise, they are usually evaluated using metrics for reliability, specificity and generalization over one or few edits. We argue that for model editing to have practical utility, we must be able to make multiple edits to the same model. With this in mind, we evaluate current model editing methods at scale, focusing on two state of the art methods - ROME and MEMIT. With the lens of scalability, we evaluate model editing methods for three crucial properties - editing proficiency, fact forgetting and downstream performance. We find that as a model is edited sequentially with multiple facts, it continually becomes less editable, forgets previously edited facts and loses the ability to perform downstream tasks. For ROME and MEMIT, this "forgetting" happens in two phases - an initial gradual but progressive forgetting phase followed by an abrupt or catastrophic forgetting. Both gradual and catastrophic forgetting limit the usefulness of model editing methods at scale - the former makes model editing less effective as multiple edits are made to the model while the latter caps the scalability of such model editing methods. Our analysis also highlights other key limitations of ROME and MEMIT at scale. With our work, we push for better evaluation of model editing and development of model editing methods keeping scalability in mind. More information can be found at the paper webpage - https://scalable-model-editing.github.io/catastrophic
@@ 1 Introduction
zsRE Who was the architect of Villa Kampen?
CounterFact Porto is a twin city of
@@ 2 Methods, Models and Datasets
In this paper, we focus on two prominent model editing methods in literature - Rank-One Memory Editing (ROME) Meng et al. (2022a) and Mass-Editing Memory in a Transformer (MEMIT) Meng et al. (2022b), whereas Model Editor Networks using Gradient Decomposition (MEND) Mitchell et al. (2021) and fine-tuning are used as baselines. MEND is a hypernetwork Chauhan et al. (2023) based model editing method that generates the weight updates for the model being edited. ROME and MEMIT first localize knowledge within a model using causal tracing Vig et al. (2020) and then update the weights of the selected layers to inject knowledge. The major difference between ROME and MEMIT is that while ROME works under the assumption that knowledge in LLMs can be updated by updating a single layer, MEMIT updates the weights of multiple layers. We evaluate these model editing methods over two models - GPT2-XL (1.5B) Radford et al. (2019) and GPT-J (6B) Wang and Komatsuzaki (2021). Note that all prior works (Mitchell et al., 2021; Meng et al., 2022a, b) edit base language models and not chat models. We refer the reader to appendix 4.1 for a detailed survey of model editing techniques. This section will be added to the main paper upon acceptance.
@@ 3.3 Downstream Evaluation of Edited Models
DISTANCE $3.339\times 10^{-4}$ $8.156\times 10^{-7}$
@@ 4 Scaling MEMIT
4.1 Related Work
In this paper, we focus on model editing methods that modify the base language model’s parameters. Some of these methods (De Cao et al., 2021; Mitchell et al., 2021) require training a hypernetwork Chauhan et al. (2023) that generates new weights for the model being edited. Other methods (Meng et al., 2022a, b; Li et al., 2023a) directly update specific parts of the model after locating stored facts inside it. Gupta et al. (2024) unify these methods under the same framework called the preservation-memorization framework and enable batched editing with ROME, an algorithm they call EMMET. Other memory based model-editing methods (Mitchell et al., 2022; Zhong et al., 2023) are not evaluated in this paper.
@@ Appendix A Appendix
A.1 zsRE Compatibility
@@
Task Few-Shot Prompt
SST-2 Review : excruciatingly unfunny and pitifully unromantic
@@
A.2 Model Editing Implementation Details
@@
A.3 Downstream Evaluation Details
@@
Task Few-Shot Prompt
MRPC Are the sentences paraphrases of each other.
@@
Task Few-Shot Prompt
COLA Is this sentence linguistically acceptable?
@@
Task Few-Shot Prompt
NLI Cyrus captured Babylon without a battle, and remedied the evils done by previous Assyrian and Babylonian rulers by sending prisoners in Babylonia back to their original homelands and aiding in the restoration of temples of the gods of various nations.
@@
A.4 Additional Scaling Experiments
A.4.1 Editing Proficiency
@@ A.5 Generation Examples
Prompt Generated Text
@@
Table 8: Text generated by GPT-J post the point of catastrophic forgetting when edited using ROME.
A.6 Background
In this section we will explain the details of four model editing algorithms explored in this paper: ROME Meng et al. (2022a), MEMIT Meng et al. (2022b), and MEND Mitchell et al. (2021), and Fine-Tuning.
A.6.1 ROME
Building off the discovery that feed-forward layers of a transformer function as key-value memories Geva et al. (2020), where neurons from $W_{fc}^{(l)}$ and $W_{proj}^{(l)}$ emulate keys and values respectively, Meng et al. (2022a) hypothesize that insertion of new knowledge can take the form of some linear transformation $W$ such that $WK\approx V$ where $K$ and $V$ are the vector of keys and values respectively. For an updated fact represented by the key-value pair $(k_{*},v_{*})$, the constrained optimization problem can be summarized as follows
$\text{min}\lVert\hat{W}K-V\rVert\ni\hat{W}k_{*}=v_{*}$ (1)
With the solution $\hat{W}=W+\Lambda(C^{-1}k_{*})^{\top}$ where $C=KK^{\top}$ and $\Lambda=\frac{v_{*}-Wk_{*}}{(C^{-1}k_{*})^{\top}k_{*}}$. The full derivation for the solution can be found in Appendix A in Meng et al. (2022a). To find the optimal $k_{*}$, inputs $x$ are taken where the subject $s$ is represented in the last token. $k_{*}$ is given by
$k_{*}=\frac{1}{N}\sum_{j=1}^{N}k(x_{j}+s)\text{, where }\\ k(x)=\sigma(W_{fc}^{(l^{*})}\gamma(a_{[x],i}^{(l^{*})}+h_{[x],i}^{(l^{*}-1)}))$ (2)
where $l^{*}$ is the desired layer, $i$ is the last subject token index, $h_{[x],i}^{(l^{*}-1)}$ is the hidden state of the previous layer, and $a_{[x],i}^{(l^{*})}$ is the global attention of the hidden layer. Here, $N$ is set to 50, since the average is taken over 50 sampled prefixes $x_{j}$. Optimal $v_{*}=\text{argmin}_{z}\mathcal{L}(z)$ where
$\mathcal{L}(z)=\frac{1}{N}\sum_{j=1}^{N}-\log(\mathbb{P}_{G(m_{i}^{(l^{*})}:=z)}[o^{*}|x_{j}+p])\\ +D_{KL}(\mathbb{P}_{G(m_{i^{\prime}}^{(l^{*})}:=z)}[x|p^{\prime}]\|\mathbb{P}_{G}[x|p^{\prime}])$ (3)
z is a vector that is substituted as the i-th token of the output to the MLP layer that enables the desired change to be realized. $G()$ substitutes the specified hidden state with the modified version. $p$ is the factual prompt, while $p^{\prime}$ is the factual prompt rewritten in a form that begins with the subject. Given these prompts, $o^{*}$ is the new object. $v_{*}$ is solved using an Adam optimizer with a learning rate of 0.5 and weight decay rate of $1.5\times 10^{-3}$. Following this, we compute the updates to the MLP weights using equation 1. ROME updates weights for GPT2-XL and GPT-J at layers 18 and 6 respectively.
A.6.2 MEMIT
Rather than overburdening one layer with an update, Meng et al. (2022b) introduces MEMIT as a means of distributing the impact of the update across multiple layers. In doing so, they are able to largely scale the number of edits they can reliably make. In order to express the update, we want to find some $z_{i}=h_{i}^{L}+\delta_{i}$ such that, when substituted in place of $h_{i}^{L}$ at layer L, it is successful. We find this by optimizing $\delta_{i}$ using
$z_{i}=h_{i}^{L}+\\ \text{argmin}_{\delta_{i}}\frac{1}{P}\sum_{j=1}^{P}-\log\mathbb{P}_{G(h_{i}^{L}+=\delta_{i})}[o_{i}|x_{j}\oplus p(s_{i},r_{i})]$ (4)
for the desired edit object $o_{i}$ and set of prompts ${x_{j}\oplus p(s_{i},r_{i})}$. Here, $x_{j}$ is a set of prefixes and $p(s_{i},r_{i})$ is a prompt generated from the edit subject $s_{i}$ and relation $r_{i}$. We want to find some update $\Delta^{l}$ for every layer $l\in R$ for a set of layers $R$ so that
$\hat{W}_{\text{out}}^{l}:=W_{\text{out}}^{l}+\Delta^{l}\text{ for all }l\in R\\ \text{such that }\text{min}_{\Delta^{l}}\sum_{i}\lVert z_{i}-\hat{h}_{i}^{L}\rVert^{2}$ (5)
where $\text{where }\hat{h}_{i}^{L}=h_{i}^{0}+\sum_{l=1}^{L}a_{i}^{l}\\ +\sum_{l=1}^{L}\hat{W}_{\text{out}}^{l}\sigma(W_{\text{in}}^{l}\gamma(h_{t}^{l-1}))$ (6)
The closed form solution to this update is given by $\Delta^{l}=R^{l}K^{l\top}(C+K^{l}K^{l\top})^{-1}$. The full derivation can be found in Meng et al. (2022b) section 4.2. To solve this, we need to find $K^{l}=[k_{1}^{l},k_{2}^{l},...,k_{n}^{l}]$ and $R^{l}=[r_{1}^{l},r_{2}^{l},...,r_{n}^{l}]$. This is found using
$k_{i}^{l}=\frac{1}{P}\sum_{j=1}^{P}k(x_{j}+s_{i})\text{, }\\ \text{where }k(x)=\sigma(W_{\text{in}}^{l}\gamma(h_{i}^{l-1}(x)))$ (7)
In this paper, Meng et al. (2022b) define $m_{[t]}^{l}=W_{\text{out}}^{l}(\sigma(W_{\text{in}}^{l}\gamma(h_{[t]}^{l-1})))$. Given this, define
$m_{i}^{l}=W_{\text{out}}k_{i}^{l}+r_{i}^{l}\\ \text{where }r_{i}^{l}=\frac{z_{i}-h_{i}^{L}}{L-l+1}$ (8)
Note that the denominator of $r_{i}^{l}$ allows us to spread out the burden across multiple layer, allowing for a more scalable algorithm. It is hard to compute $C^{l}$ exactly, however it can be reliably estimated using $C^{l}=\lambda\mathbb{E}_{k}[k^{l}k^{l\top}]$ over randomly sampled inputs, where $\lambda=1.5\times 10^{4}$. Incorporating the update gives us our desired new weights $\hat{W}_{\text{out}}^{l}$
A.6.3 MEND
Using the fact that the gradient of loss L with respect to the weights $W_{\ell}$ of layer $\ell$ of an MLP has a rank-one decomposition such that $\nabla_{W_{\ell}}L=\sum_{i=1}^{B}\delta_{\ell+1}^{i}u_{\ell}^{i\top}$ for a batch $B$, Mitchell et al. (2021) are able to construct an editor network $g_{\ell}$ to generate the weight updates. Here, $\delta_{\ell+1}^{i}$ is the gradient for element $i$ for the preactivations of layer $\ell+1$ and $u_{\ell}^{i}$ are the inputs of element $i$ into layer $\ell$.
To characterize these updates, MEND employs functions that map $\delta_{\ell+1}^{i}$ and $u_{\ell}^{i}$ to a pseudo-decomposition $\tilde{\delta}_{\ell+1}^{i}$ and $\tilde{u}_{\ell}^{i}$ such that ~$\tilde{\nabla}_{W_{\ell}}L=\sum_{i=1}^{B}\tilde{\delta}_{\ell+1}^{i}\tilde{u}_{\ell}^{i\top}$. Letting $z_{\ell}=\text{concat}(\delta_{\ell+1}\text{, }u_{\ell})$, the form of the network is
$h_{\ell}=z_{\ell}+\sigma(s_{\ell}^{1}\odot(U_{1}V_{1}z_{\ell}+b)+o_{\ell}^{1})$ (9)
$g(z_{\ell})=h_{\ell}+\sigma(s_{\ell}^{2}\odot U_{2}V_{2}h_{\ell}+o_{\ell}^{2})$ (10)
where $\sigma$ is the ReLu activation function and $U_{j}$, $V_{j}$ are a low rank decomposition of MEND’s weight for layer $j$. Note that, because of the difference in dimensions between weight matrices across layers, MEND learns different parameters for each unique shape of weight matrices to be edited. Additionally, layer-wise offset and scale parameters $o_{\ell}$ and $s_{\ell}$ are learned for both $h_{\ell}$ and $g_{\ell}$. The final update is given by $\tilde{W}=W-\alpha_{\ell}\tilde{\nabla}_{W_{\ell}}$ with $\alpha_{\ell}$ being another learned parameter per layer.
Given the original weights $W$ and the updated weights $\tilde{W}$, loss is computed aggregating two training losses, editing success and locality. For a desired edit $(x_{e},y_{e})$, $(x_{e}^{\prime},y_{e}^{\prime})$ is defined as a semantically equivalent wording of the edit. Editing loss is defined as $L_{e}=-\log_{p_{\theta_{\tilde{W}}}}p(y_{e}^{\prime}|x_{e}^{\prime})$. $x_{\text{loc}}$ is defined as a locality sample, which is randomly sampled to test the edited model’s impact on information unrelated to the edit. The corresponding locality loss is $L_{\text{loc}}=D_{KL}(p_{\theta_{W}}(\cdot|x_{\text{loc}})\|p_{\theta_{\tilde{W}}}(\cdot|x_{\text{loc}}))$.
The total loss is computed as $L_{\text{MEND}}=c_{e}L_{e}(\theta_{\tilde{W}})+L_{\text{loc}}(\theta_{W},\theta_{\tilde{W}})$ where $c_{e}=0.1$. The total loss is optimized using the Adam optimizer. For GPT2-XL, we edit layers 46, 47, and 48. For GPTJ, we edit layers 26, 27, and 28.
A.6.4 FT
The Fine-Tuning procedures used in this paper follow from Meng et al. (2022b) and Meng et al. (2022a)’s implementation for both GPT-J and GPT2-XL. MLP weights for a single layer are fine-tuned for both models. We use a constrained fine tuning approach where we add a $L_{\infty}$ constraint such that $\lVert\theta_{G}-\theta_{G^{\prime}}\rVert_{\infty}\leq\epsilon$ at each gradient step. For the constraint, $\epsilon=5e-4$ for GPT2-XL and $\epsilon=5e-5$ for GPT-J. It is optimized using Adam with a learning rate of 5e-4 for both GPT2-XL and GPT-J. We fine tune layers 18 and 6 for GPT2-XL and GPT-J respectively.